Doctor’s Order: Nafcillin 500 mg po pc; Available: Nafcillin 1 gm tab (scored). How many tab will you administer per day? To nearest.



The baby is getting 5ml/hr of D10W. How many grams of dextrose will they receive for the whole day?

PLACE NUMBER ONLY IN LINE PROVIDED
Answer #1

1. Number of tabs administered per day = 1.5 tabs

Calculation:

Nafcillin 1 gm = 1000 mg

number of 500 mg tabs to be administered per day = 3

Three 500 mg tabs = one-and-a-half 1000 mg tabs per day

2. Dextrose for one whole day = 12 grams

D10W given = 5 mL/ hr

D10W given in 24 hours = 5*24 = 120 mL

Dextrose in 100 mL D10W = 10 grams

Dextrose in 120 mL D10W = 12 grams